# Edwards-Wilkinson limit for a stochastic advection-diffusion PDE

### v1

@@ 2.1. Correlation Functions and a priori estimates
$\big{\|}P_{t-r}\nabla\cdot(\theta_{r}\sigma_{k})\big{\|}_{\dot{H}^{-\alpha}}^{2}$ $=\int_{\mathbb{R}^{d}}|\xi|^{-2\alpha}\big{|}\mathcal{F}\big{(}P_{t-r}\nabla\cdot(\theta_{r}\sigma_{k})\big{)}(\xi)\big{|}^{2}\,\mathrm{d}\xi$
$=\int_{\mathbb{R}^{d}}|\xi|^{-2\alpha}{\rm e}^{-2(\kappa+\nu)|\xi|^{2}(t-r)}\big{|}\xi\cdot\mathcal{F}(\theta_{r}\sigma_{k})(\xi)\big{|}^{2}\,\mathrm{d}\xi$
$=(2\pi)^{-d}\int_{\mathbb{R}^{d}}|\xi|^{-2\alpha}{\rm e}^{-2(\kappa+\nu)|\xi|^{2}(t-r)}\big{|}\xi\cdot(\widehat{\theta}_{r}\ast\widehat{\sigma}_{k})(\xi)\big{|}^{2}\,\mathrm{d}\xi,$
@@
$\big{[}\mathbb{E}\|Z_{t}\|_{\dot{H}^{-\alpha}}^{2q}\big{]}^{1/q}$ $\lesssim_{d,q}\bigg{[}\mathbb{E}\Big{(}\sum_{k}\int_{0}^{t}\!\int_{\mathbb{R}^{d}}|\xi|^{-2\alpha}{\rm e}^{-2(\kappa+\nu)|\xi|^{2}(t-r)}\big{|}\xi\cdot(\widehat{\theta}_{r}\ast\widehat{\sigma}_{k})(\xi)\big{|}^{2}\,\mathrm{d}\xi\,\mathrm{d}r\Big{)}^{q}\bigg{]}^{1/q}$
$=\bigg{[}\mathbb{E}\Big{(}\int_{\mathbb{R}^{d}}|\xi|^{-2\alpha}\int_{0}^{t}{\rm e}^{-2(\kappa+\nu)|\xi|^{2}(t-r)}\sum_{k}\big{|}\xi\cdot(\widehat{\theta}_{r}\ast\widehat{\sigma}_{k})(\xi)\big{|}^{2}\,\mathrm{d}r\,\mathrm{d}\xi\Big{)}^{q}\bigg{]}^{1/q}.$
By Proposition 1.4, we see that
(2.12) $\sum_{k}|\xi\cdot(\widehat{\theta}_{r}\ast\widehat{\sigma}_{k})(\xi)|^{2}$
$=\lim_{N\to\infty}\sum_{k=0}^{N}\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\widehat{\theta}_{r}(\xi-\eta)~\xi^{T}\widehat{\sigma}_{k}(\eta)\overline{\widehat{\sigma}_{k}(\zeta)^{T}}\xi~\overline{\widehat{\theta}_{r}(\xi-\zeta)}\,\mathrm{d}\eta\,\mathrm{d}\zeta$
$=\int_{\mathbb{R}^{d}}\xi^{T}\widehat{Q}(\eta)\xi\,|\widehat{\theta}_{r}(\xi-\eta)|^{2}\,\mathrm{d}\eta.$
@@
$\big{[}\mathbb{E}\|Z_{t}\|_{\dot{H}^{-\alpha}}^{2q}\big{]}^{1/q}\lesssim\bigg{[}\mathbb{E}\Big{(}\int_{\mathbb{R}^{d}}\!|\xi|^{-2\alpha}\!\int_{0}^{t}\!{\rm e}^{-2(\kappa+\nu)|\xi|^{2}(t-r)}\xi^{T}\big{(}|\widehat{\theta}_{r}|^{2}\ast\widehat{Q}\big{)}(\xi)\xi\,\mathrm{d}r\,\mathrm{d}\xi\Big{)}^{q}\bigg{]}^{1/q}.$
@@
$\big{|}\big{(}|\widehat{\theta}_{r}|^{2}\ast\widehat{Q}\big{)}(\xi)\big{|}\leqslant\big{\|}|\widehat{\theta}_{r}|^{2}\big{\|}_{1}\big{\|}\widehat{Q}\big{\|}_{\infty}=\|\widehat{\theta}_{r}\|_{2}^{2}\big{\|}\widehat{Q}\big{\|}_{\infty}=\|\theta_{r}\|_{2}^{2}\big{\|}\widehat{Q}\big{\|}_{\infty}$.
Plugging this to the previous bound yields
$\mathbb{E}\Big{[}\|Z_{t}\|_{\dot{H}^{-\alpha}_{x}}^{2q}\Big{]}^{1/q}\lesssim\bigg{[}\mathbb{E}\Big{(}\int_{0}^{t}\!\!\int_{\mathbb{R}^{d}}\!|\xi|^{-2\alpha+2}{\rm e}^{-2(\kappa+\nu)|\xi|^{2}(t-r)}\|\theta_{r}\|_{2}^{2}\big{\|}\widehat{Q}\big{\|}_{\infty}\,\mathrm{d}r\,\mathrm{d}\xi\Big{)}^{q}\bigg{]}^{1/q}.$
@@
$\mathbb{E}\Big{[}\|\theta_{t}-\bar{\theta}_{t}\|_{\dot{H}^{-\alpha}_{x}}^{2q}\Big{]}^{1/q}\lesssim\big{\|}\widehat{Q}\big{\|}_{\infty}\int_{0}^{t}\!\!\int_{\mathbb{R}^{d}}\!|\xi|^{-2\alpha+2}{\rm e}^{-2(\kappa+\nu)|\xi|^{2}(t-r)}\mathbb{E}\big{[}\|\theta_{r}\|_{2}^{2q}\big{]}^{1/q}\,\mathrm{d}r\,\mathrm{d}\xi.$
@@
$\big{[}\mathbb{E}\|Z_{t}\|_{\dot{H}^{-\alpha}}^{2q}\big{]}^{1/q}$ $\lesssim\|\varphi\|_{L^{2}_{x}}^{2}\big{\|}\widehat{Q}\big{\|}_{\infty}\int_{0}^{t}\!\!\int_{\mathbb{R}^{d}}\!|\xi|^{-2\alpha+2}{\rm e}^{-2(\kappa+\nu)|\xi|^{2}(t-r)}\,\mathrm{d}r\,\mathrm{d}\xi$
$=\|\varphi\|_{L^{2}_{x}}^{2}\big{\|}\widehat{Q}\big{\|}_{\infty}\int_{\mathbb{R}^{d}}|\xi|^{-2\alpha}\frac{1-{\rm e}^{-2(\kappa+\nu)|\xi|^{2}t}}{2(\kappa+\nu)}\,\mathrm{d}\xi$
$\leqslant\|\varphi\|_{L^{2}_{x}}^{2}\big{\|}\widehat{Q}\big{\|}_{\infty}\int_{\mathbb{R}^{d}}|\xi|^{-2\alpha}\big{[}(2\nu)^{-1}\wedge(|\xi|^{2}t)\big{]}\,\mathrm{d}\xi.$
@@
$\big{[}\mathbb{E}\|Z_{t}\|_{\dot{H}^{-\alpha}}^{q}\big{]}^{1/q}\leqslant\big{[}\mathbb{E}\|Z_{t}\|_{\dot{H}^{-\alpha}}^{2q}\big{]}^{1/2q}\lesssim_{q,d,\alpha}\|\varphi\|_{L^{2}_{x}}\big{\|}\widehat{Q}\big{\|}_{\infty}^{1/2}t^{(2\alpha-d)/4}.$
It remains to show that the above estimate can be improved by inserting $\sup_{t\in[0,T]}$ in the expectation. This can be done in the same way as the end of proof of [23, Lemma 3.1]. Hence we omit the details here. ∎
2.2. Proof of Theorem 1.9: The compressible case
Here, we prove Theorem 1.9, under the Assumption 1.2. As mentioned in the introduction, the idea is the following: First, we will prove that $(\mathcal{X}_{n})_{n}$, defined in (1.4), is tight and we will characterize the law of all limiting points.
Recall $\theta_{n}(t,x)=n^{d}\theta(n^{2}t,nx)$, which satisfies the equation
(2.13) $\partial_{t}\theta_{n}+\nabla\cdot(\theta_{n}V^{n})=(\kappa+\nu)\Delta\theta_{n},$
with the rescaled initial condition $\theta(0,x)=n^{-d}\varphi(n^{-1}x)$. Therefore, the initial condition for (2.13) is
$\theta_{n}(0,x)=n^{d}\theta(0,nx)=\varphi(x)$
which is independent of $n$.
The noise $V^{n}$ in (2.13) has covariance function $Q^{n}(x)=Q(nx),\,x\in\mathbb{R}^{d}$; therefore, it has the Fourier transform
(2.14) $\widehat{Q^{n}}(\xi)=n^{-d}\widehat{Q}(n^{-1}\xi).$
Note that $V^{n}=\dot{W}^{n}$ and thus by (1.6),
as a result,
by the scaling property of Brownian motions. Finally, note that $Q^{n}(0)=Q(0)=2\nu I_{d\times d}$ is independent of $n$.
By the above discussions, for $\alpha\in\big{(}\frac{d}{2},\frac{d}{2}+1\big{)}$, we can apply Proposition 2.6 to (2.13) and obtain
(2.16) $\mathbb{E}\Big{[}\sup_{t\in[0,T]}\|\theta_{n}(t)-\bar{\theta}(t)\|_{\dot{H}^{-\alpha}_{x}}^{q}\Big{]}^{1/q}$ $\lesssim_{d,\alpha,q,T}\|\varphi\|_{2}\big{\|}\widehat{Q^{n}}\big{\|}_{\infty}^{1/2}$
$=\|\varphi\|_{2}\big{\|}\widehat{Q}\big{\|}_{\infty}^{1/2}n^{-d/2},$
where in the second step we have used the equality in (2.14). In particular, taking $\alpha=\frac{d}{2}+\varepsilon$ with $\varepsilon\in(0,1)$, we arrive at
$\mathbb{E}\Big{[}\sup_{t\in[0,T]}\|\theta_{n}(t)-\bar{\theta}(t)\|_{\dot{H}^{-d/2-\varepsilon}_{x}}^{q}\Big{]}^{1/q}$ $\lesssim_{d,\alpha,q,T}\|\varphi\|_{2}\big{\|}\widehat{Q}\big{\|}_{\infty}^{1/2}n^{-d/2}.$
as a result,
(2.17) $\mathbb{E}\Big{[}\sup_{t\in[0,T]}\big{\|}n^{d/2}(\theta_{n}(t)-\bar{\theta}(t))\big{\|}_{\dot{H}^{-d/2-\varepsilon}_{x}}^{q}\Big{]}^{1/q}$ $\lesssim_{d,\alpha,q,T}\|\varphi\|_{2}\big{\|}\widehat{Q}\big{\|}_{\infty}^{1/2}.$
Recalling (1.4), we have proved
$\mathbb{E}\Big{[}\sup_{t\in[0,T]}\|\mathcal{X}_{n}(t)\|_{\dot{H}^{-d/2-\varepsilon}_{x}}^{q}\Big{]}^{1/q}\lesssim_{d,\alpha,q}\|\varphi\|_{2}\big{\|}\widehat{Q}\big{\|}_{\infty}^{1/2}.$
Next, note that $\mathcal{X}_{n}$ has the expression
(2.18) $\mathcal{X}_{n}(t)$ $=n^{d/2}\int_{0}^{t}P_{t-r}\nabla\cdot(\theta_{n}(r)V^{n}(\,\mathrm{d}r))$
$=n^{d/2}\sum_{k}\int_{0}^{t}P_{t-r}\nabla\cdot(\theta_{n}(r)\sigma^{n}_{k})\,\mathrm{d}B^{n}_{k}(r),$
@@
$\mathcal{X}_{n}(t)-\mathcal{X}_{n}(s)$ $=n^{d/2}\int_{0}^{s}\sum_{k}(P_{t-r}-P_{s-r})(\nabla\cdot(\sigma^{n}_{k}\theta_{n}(r)))\,\mathrm{d}B^{n}_{k}(r)$
$\quad+n^{d/2}\int_{s}^{t}\sum_{k}P_{t-r}(\nabla\cdot(\sigma^{n}_{k}\theta_{n}(r)))\,\mathrm{d}B^{n}_{k}(r)$
@@
$\mathbb{E}\|I^{n}_{1}\|_{H^{-\alpha}}^{q}$ $\lesssim_{q}n^{qd/2}\mathbb{E}\bigg{(}\int_{0}^{s}\sum_{k}\big{\|}(P_{t-s}-I)P_{s-r}(\nabla\cdot(\sigma^{n}_{k}\theta_{n}(r)))\big{\|}_{H^{-\alpha}}^{2}\,\mathrm{d}r\bigg{)}^{q/2}$
$\lesssim n^{qd/2}\mathbb{E}\bigg{(}\int_{0}^{s}\sum_{k}|t-s|^{\delta}\big{\|}P_{s-r}(\nabla\cdot(\sigma^{n}_{k}\theta_{n}(r)))\big{\|}_{H^{-\alpha+\delta}}^{2}\,\mathrm{d}r\bigg{)}^{q/2}$
$\lesssim_{\nu}(n^{d}|t-s|^{\delta})^{q/2}\mathbb{E}\bigg{(}\int_{0}^{s}\sum_{k}\frac{1}{|s-r|^{1-\varepsilon}}\big{\|}\nabla\cdot(\sigma^{n}_{k}\theta_{n}(r))\big{\|}_{H^{-\alpha+\delta-1+\varepsilon}}^{2}\,\mathrm{d}r\bigg{)}^{q/2}$
$\lesssim(n^{d}|t-s|^{\delta})^{q/2}\mathbb{E}\bigg{(}\int_{0}^{s}\frac{1}{|s-r|^{1-\varepsilon}}\sum_{k}\big{\|}\sigma^{n}_{k}\theta_{n}(r)\big{\|}_{H^{-\alpha+\delta+\varepsilon}}^{2}\,\mathrm{d}r\bigg{)}^{q/2},$
where $\varepsilon$ is small enough and in the third step we have used semigroup property. We have
$\sum_{k}\big{\|}\sigma^{n}_{k}\theta_{n}(r)\big{\|}_{H^{-\alpha+\delta+\varepsilon}}^{2}$ $=\sum_{k}\int_{\mathbb{R}^{d}}\langle\xi\rangle^{-2(\alpha-\delta-\varepsilon)}\big{|}\mathcal{F}(\theta_{n}(r)\sigma^{n}_{k})(\xi)\big{|}^{2}\,\mathrm{d}\xi$
$=\frac{1}{(2\pi)^{d}}\int_{\mathbb{R}^{d}}\langle\xi\rangle^{-2(\alpha-\delta-\varepsilon)}\sum_{k}\big{|}\big{(}\widehat{\theta}_{n}(r)\ast\widehat{\sigma^{n}_{k}}\big{)}(\xi)\big{|}^{2}\,\mathrm{d}\xi.$
@@
$\sum_{k}\big{|}\big{(}\widehat{\theta}_{n}(r)\ast\widehat{\sigma^{n}_{k}}\big{)}(\xi)\big{|}^{2}$ $=\sum_{k}\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\widehat{\theta}_{n}(r,\xi-\eta)\overline{\widehat{\theta}_{n}(r,\xi-\zeta)}\widehat{\sigma^{n}_{k}}(\eta)\cdot\overline{\widehat{\sigma^{n}_{k}}(\zeta)}\,\mathrm{d}\eta\,\mathrm{d}\zeta$
@@
$\sum_{k}\big{\|}\sigma^{n}_{k}\theta_{n}(r)\big{\|}_{H^{-\alpha+\delta+\varepsilon}}^{2}$ $=\frac{1}{(2\pi)^{d}}\int_{\mathbb{R}^{d}}\langle\xi\rangle^{-2(\alpha-\delta-\varepsilon)}\Big{(}\big{|}\widehat{\theta}_{n}(r)\big{|}^{2}\ast{\rm Tr}\big{(}\widehat{Q^{n}}\big{)}\Big{)}(\xi)\,\mathrm{d}\xi.$
@@
$\Big{(}\big{|}\widehat{\theta}_{n}(r)\big{|}^{2}\ast{\rm Tr}\big{(}\widehat{Q^{n}}\big{)}\Big{)}(\xi)\leqslant\Big{\|}\big{|}\widehat{\theta}_{n}(r)\big{|}^{2}\Big{\|}_{1}\big{\|}{\rm Tr}\big{(}\widehat{Q^{n}}\big{)}\big{\|}_{\infty}\leqslant\big{\|}\widehat{\theta}_{n}(r)\big{\|}_{2}^{2}\big{\|}\widehat{Q^{n}}\big{\|}_{\infty}$.
@@
$\Big{(}\big{|}\widehat{\theta}_{n}(r)\big{|}^{2}\ast{\rm Tr}\big{(}\widehat{Q^{n}}\big{)}\Big{)}(\xi)\leqslant\big{\|}\widehat{\theta}_{n}(r)\big{\|}_{2}^{2}\,n^{-d}\big{\|}\widehat{Q}\big{\|}_{\infty}=n^{-d}\big{\|}\widehat{Q}\big{\|}_{\infty}\|\theta_{n}(r)\|_{2}^{2}.$
@@
$\sum_{k}\big{\|}\sigma^{n}_{k}\theta_{n}(r)\big{\|}_{H^{-\alpha+\delta+\varepsilon}}^{2}$ $\leqslant n^{-d}\big{\|}\widehat{Q}\big{\|}_{\infty}\|\theta_{n}(r)\|_{2}^{2}\int_{\mathbb{R}^{d}}\langle\xi\rangle^{-2(\alpha-\delta-\varepsilon)}\,\mathrm{d}\xi$
$\lesssim n^{-d}\big{\|}\widehat{Q}\big{\|}_{\infty}\|\theta_{n}(r)\|_{2}^{2},$

### v2

@@ 2.1. Correlation Functions and a priori estimates
$\big\|P_{t-r}\nabla\cdot(\theta_{r}^{n}\sigma_{k}^{n})\big\|_{\dot{H}^{-\alpha}}^{2}$ $=\int_{\mathbb{R}^{d}}|\xi|^{-2\alpha}\big|\mathcal{F}\big(P_{t-r}\nabla\cdot(\theta_{r}^{n}\sigma_{k}^{n})\big)(\xi)\big|^{2}\,\mathrm{d}\xi$
$=\int_{\mathbb{R}^{d}}|\xi|^{-2\alpha}{\rm e}^{-2(\kappa+\nu)|\xi|^{2}(t-r)}\big|\xi\cdot\mathcal{F}(\theta_{r}^{n}\sigma_{k}^{n})(\xi)\big|^{2}\,\mathrm{d}\xi$
$=(2\pi)^{-d}\int_{\mathbb{R}^{d}}|\xi|^{-2\alpha}{\rm e}^{-2(\kappa+\nu)|\xi|^{2}(t-r)}\big|\xi\cdot\big(\widehat{\theta^{n}}(r)\ast\widehat{\sigma^{n}_{k}}\big)(\xi)\big|^{2}\,\mathrm{d}\xi,$
@@
$\big[\mathbb{E}\|Z_{t}^{n}\|_{\dot{H}^{-\alpha}}^{2q}\big]^{1/q}$
$\lesssim_{d,q}\bigg[\mathbb{E}\Big(\sum_{k}\int_{0}^{t}\!\int_{\mathbb{R}^{d}}|\xi|^{-2\alpha}{\rm e}^{-2(\kappa+\nu)|\xi|^{2}(t-r)}\big|\xi\cdot\big(\widehat{\theta^{n}}(r)\ast\widehat{\sigma^{n}_{k}}\big)(\xi)\big|^{2}\,\mathrm{d}\xi\,\mathrm{d}r\Big)^{q}\bigg]^{1/q}$
$=\bigg[\mathbb{E}\Big(\int_{\mathbb{R}^{d}}|\xi|^{-2\alpha}\int_{0}^{t}{\rm e}^{-2(\kappa+\nu)|\xi|^{2}(t-r)}\sum_{k}\big|\xi\cdot\big(\widehat{\theta^{n}}(r)\ast\widehat{\sigma^{n}_{k}}\big)(\xi)\big|^{2}\,\mathrm{d}r\,\mathrm{d}\xi\Big)^{q}\bigg]^{1/q}.$
By Proposition 1.4, and the scaling properties of the Fourier transform, we see that
(2.15) $\sum_{k}\big|\xi\cdot(\widehat{\theta^{n}}(r)\ast\widehat{\sigma^{n}_{k}})(\xi)\big|^{2}$
$=\lim_{N\to\infty}\sum_{k=0}^{N}\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\widehat{\theta^{n}}(r,\xi-\eta)~\xi^{\ast}\widehat{\sigma^{n}_{k}}(\eta)\overline{\widehat{\sigma^{n}_{k}}(\zeta)^{\ast}}\xi~\overline{\widehat{\theta^{n}}(r,\xi-\zeta)}\,\mathrm{d}\eta\,\mathrm{d}\zeta$
$=\int_{\mathbb{R}^{d}}\xi^{\ast}\widehat{Q^{n}}(\eta)\xi\,\big|\widehat{\theta^{n}}(r,\xi-\eta)\big|^{2}\,\mathrm{d}\eta.$
@@
$\big[\mathbb{E}\|Z_{t}^{n}\|_{\dot{H}^{-\alpha}}^{2q}\big]^{1/q}\lesssim\bigg[\mathbb{E}\Big(\int_{\mathbb{R}^{d}}\!|\xi|^{-2\alpha}\!\int_{0}^{t}\!{\rm e}^{-2(\kappa+\nu)|\xi|^{2}(t-r)}\xi^{\ast}\big(|\widehat{\theta^{n}}(r)|^{2}\ast\widehat{Q^{n}}\big)(\xi)\xi\,\mathrm{d}r\,\mathrm{d}\xi\Big)^{q}\bigg]^{1/q}.$
@@
$\big|\big(|\widehat{\theta^{n}}(r)|^{2}\ast\widehat{Q^{n}}\big)(\xi)\big|\leqslant\big\||\widehat{\theta^{n}}(r)|^{2}\big\|_{1}\big\|\widehat{Q^{n}}\big\|_{\infty}=\|\widehat{\theta^{n}}(r)\|_{2}^{2}\big\|\widehat{Q^{n}}\big\|_{\infty}=\|\theta^{n}_{r}\|_{2}^{2}\big\|\widehat{Q^{n}}\big\|_{\infty}$.
Plugging this into the previous bound yields
$\big[\mathbb{E}\|Z_{t}^{n}\|_{\dot{H}^{-\alpha}}^{2q}\big]^{1/q}\lesssim\bigg[\mathbb{E}\Big(\int_{0}^{t}\!\!\int_{\mathbb{R}^{d}}\!|\xi|^{-2\alpha+2}{\rm e}^{-2(\kappa+\nu)|\xi|^{2}(t-r)}\|\theta^{n}_{r}\|_{2}^{2}\big\|\widehat{Q^{n}}\big\|_{\infty}\,\mathrm{d}r\,\mathrm{d}\xi\Big)^{q}\bigg]^{1/q}.$
@@
$\big[\mathbb{E}\|Z_{t}^{n}\|_{\dot{H}^{-\alpha}}^{2q}\big]^{1/q}\lesssim\big\|\widehat{Q^{n}}\big\|_{\infty}\int_{0}^{t}\!\!\int_{\mathbb{R}^{d}}\!|\xi|^{-2\alpha+2}{\rm e}^{-2(\kappa+\nu)|\xi|^{2}(t-r)}\mathbb{E}\big[\|\theta^{n}_{r}\|_{2}^{2q}\big]^{1/q}\,\mathrm{d}r\,\mathrm{d}\xi.$
@@
$\big[\mathbb{E}\|Z_{t}^{n}\|_{\dot{H}^{-\alpha}}^{2q}\big]^{1/q}$ $\lesssim\|\varphi\|_{L^{2}_{x}}^{2}\big\|\widehat{Q^{n}}\big\|_{\infty}\int_{0}^{t}\!\!\int_{\mathbb{R}^{d}}\!|\xi|^{-2\alpha+2}{\rm e}^{-2(\kappa+\nu)|\xi|^{2}(t-r)}\,\mathrm{d}r\,\mathrm{d}\xi$
$=\|\varphi\|_{L^{2}_{x}}^{2}\big\|\widehat{Q^{n}}\big\|_{\infty}\int_{\mathbb{R}^{d}}|\xi|^{-2\alpha}\frac{1-{\rm e}^{-2(\kappa+\nu)|\xi|^{2}t}}{2(\kappa+\nu)}\,\mathrm{d}\xi$
$\leqslant\|\varphi\|_{L^{2}_{x}}^{2}\big\|\widehat{Q^{n}}\big\|_{\infty}\int_{\mathbb{R}^{d}}|\xi|^{-2\alpha}\big[(2\nu)^{-1}\wedge(|\xi|^{2}t)\big]\,\mathrm{d}\xi.$
@@
$\big[\mathbb{E}\|Z_{t}^{n}\|_{\dot{H}^{-\alpha}}^{q}\big]^{1/q}\leqslant\big[\mathbb{E}\|Z_{t}^{n}\|_{\dot{H}^{-\alpha}}^{2q}\big]^{1/2q}\lesssim_{q,d,\alpha}\|\theta_{0}\|_{L^{2}_{x}}\big\|\widehat{Q}\big\|_{\infty}^{1/2}t^{(2\alpha-d)/4}.$
It remains to show that the above estimate can be improved by inserting $\sup_{t\in[0,T]}$ in the expectation. This can be done in the same way as the end of proof of [33, Lemma 3.1]. Hence, we omit the details here. ∎
Remark 2.8.
In the case where we do not rescale the solution and the initial data, we can show the following estimate
where $\theta_{t}$ solves (1.1) with $\varphi$ as the initial data. Indeed, we can follow the exact same computations, with the difference that we use the bound in Remark 2.5 instead of Lemma 2.4. Notice that (2.16) has the same form as (2.13), the only difference coming from the rescaling of the correlation function of the noise.
As mentioned before, estimate (2.16) is very similar to the ones in [33, Theorem 1.5], with the major technical difference being that we do not assume that $Q$ is divergence-free. Nevertheless, a similar bound to (2.16) holds under Assumption 1.1 as well. This can be seen by following the same arguments as in the proof of (2.16) but using (1.8) in place of Lemma 2.4. In fact, more generally, under Assumption 1.1 we have
(2.17) $\mathbb{E}\Big[\sup_{t\in[0,T]}\|\theta_{t}-\bar{\theta}_{t}\|_{\dot{H}^{-\alpha}_{x}}^{q}\Big]^{1/q}\lesssim_{d,\alpha,\nu,q,T}\|\theta_{0}\|_{p}\big\|\widehat{Q}\big\|_{p/(2-p)}^{1/2},$
for $p\in(1,2]$, $q\geqslant 1$ and $\alpha\in(\frac{d}{2},\frac{d}{2}+1)$. The proof (2.17) is similar to the proof of Proposition 2.7, and since we are not going to use this bound, we skip the details, cf. [33, Theorem 1.5].
Now, note that $\mathcal{X}_{n}$ has the expression
(2.18) $\mathcal{X}_{n}(t)$ $=n^{d/2}\int_{0}^{t}P_{t-r}\nabla\cdot(\theta^{n}_{r}V^{n}(\,\mathrm{d}r))$
$=n^{d/2}\sum_{k}\int_{0}^{t}P_{t-r}\nabla\cdot(\theta^{n}_{r}\sigma^{n}_{k})\,\mathrm{d}B^{n}_{k}(r),$
@@
$\mathcal{X}_{n}(t)-\mathcal{X}_{n}(s)$ $=n^{d/2}\int_{0}^{s}\sum_{k}(P_{t-r}-P_{s-r})(\nabla\cdot(\sigma^{n}_{k}\theta^{n}_{r}))\,\mathrm{d}B^{n}_{k}(r)$
$\quad+n^{d/2}\int_{s}^{t}\sum_{k}P_{t-r}(\nabla\cdot(\sigma^{n}_{k}\theta^{n}_{r}))\,\mathrm{d}B^{n}_{k}(r)$
@@
$\mathbb{E}\|I^{n}_{1}\|_{H^{-\alpha}}^{q}$ $\lesssim_{q}n^{qd/2}\mathbb{E}\bigg(\int_{0}^{s}\sum_{k}\big\|(P_{t-s}-I)P_{s-r}(\nabla\cdot(\sigma^{n}_{k}\theta^{n}_{r}))\big\|_{H^{-\alpha}}^{2}\,\mathrm{d}r\bigg)^{q/2}$
$\lesssim n^{qd/2}\mathbb{E}\bigg(\int_{0}^{s}\sum_{k}|t-s|^{\delta}\big\|P_{s-r}(\nabla\cdot(\sigma^{n}_{k}\theta^{n}_{r}))\big\|_{H^{-\alpha+\delta}}^{2}\,\mathrm{d}r\bigg)^{q/2}$
$\lesssim_{\nu}(n^{d}|t-s|^{\delta})^{q/2}\mathbb{E}\bigg(\int_{0}^{s}\sum_{k}\frac{1}{|s-r|^{1-\varepsilon}}\big\|\nabla\cdot(\sigma^{n}_{k}\theta^{n}_{r})\big\|_{H^{-\alpha+\delta-1+\varepsilon}}^{2}\,\mathrm{d}r\bigg)^{q/2}$
$\lesssim(n^{d}|t-s|^{\delta})^{q/2}\mathbb{E}\bigg(\int_{0}^{s}\frac{1}{|s-r|^{1-\varepsilon}}\sum_{k}\big\|\sigma^{n}_{k}\theta^{n}_{r}\big\|_{H^{-\alpha+\delta+\varepsilon}}^{2}\,\mathrm{d}r\bigg)^{q/2},$
where $\varepsilon$ is small enough, and in the third step we have used the semigroup property. We have
$\sum_{k}\big\|\sigma^{n}_{k}\theta^{n}_{r}\big\|_{H^{-\alpha+\delta+\varepsilon}}^{2}$ $=\sum_{k}\int_{\mathbb{R}^{d}}\langle\xi\rangle^{-2(\alpha-\delta-\varepsilon)}\big|\mathcal{F}(\theta^{n}_{r}\sigma^{n}_{k})(\xi)\big|^{2}\,\mathrm{d}\xi$
$=\frac{1}{(2\pi)^{d}}\int_{\mathbb{R}^{d}}\langle\xi\rangle^{-2(\alpha-\delta-\varepsilon)}\sum_{k}\big|\big(\widehat{\theta^{n}}(r)\ast\widehat{\sigma^{n}_{k}}\big)(\xi)\big|^{2}\,\mathrm{d}\xi.$
@@
$\sum_{k}\big|\big(\widehat{\theta^{n}}(r)\ast\widehat{\sigma^{n}_{k}}\big)(\xi)\big|^{2}$ $=\sum_{k}\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\widehat{\theta}_{n}(r,\xi-\eta)\overline{\widehat{\theta}_{n}(r,\xi-\zeta)}\widehat{\sigma^{n}_{k}}(\eta)\cdot\overline{\widehat{\sigma^{n}_{k}}(\zeta)}\,\mathrm{d}\eta\,\mathrm{d}\zeta$
@@
$\sum_{k}\big\|\sigma^{n}_{k}\theta^{n}_{r}\big\|_{H^{-\alpha+\delta+\varepsilon}}^{2}$ $=\frac{1}{(2\pi)^{d}}\int_{\mathbb{R}^{d}}\langle\xi\rangle^{-2(\alpha-\delta-\varepsilon)}\Big(\big|\widehat{\theta^{n}}(r)\big|^{2}\ast{\rm Tr}\big(\widehat{Q^{n}}\big)\Big)(\xi)\,\mathrm{d}\xi.$
@@
$\Big(\big|\widehat{\theta^{n}}(r)\big|^{2}\ast{\rm Tr}\big(\widehat{Q^{n}}\big)\Big)(\xi)\leqslant\Big\|\big|\widehat{\theta^{n}}(r)\big|^{2}\Big\|_{1}\big\|{\rm Tr}\big(\widehat{Q^{n}}\big)\big\|_{\infty}\leqslant\big\|\widehat{\theta^{n}}(r)\big\|_{2}^{2}\big\|\widehat{Q^{n}}\big\|_{\infty}$.
@@
$\Big(\big|\widehat{\theta^{n}}(r)\big|^{2}\ast{\rm Tr}\big(\widehat{Q^{n}}\big)\Big)(\xi)\leqslant\big\|\widehat{\theta^{n}}(r)\big\|_{2}^{2}\,n^{-d}\big\|\widehat{Q}\big\|_{\infty}=n^{-d}\big\|\widehat{Q}\big\|_{\infty}\|\theta^{n}_{r}\|_{2}^{2}.$
@@
$\sum_{k}\big\|\sigma^{n}_{k}\theta^{n}_{r}\big\|_{H^{-\alpha+\delta+\varepsilon}}^{2}$ $\leqslant n^{-d}\big\|\widehat{Q}\big\|_{\infty}\|\theta^{n}_{r}\|_{2}^{2}\int_{\mathbb{R}^{d}}\langle\xi\rangle^{-2(\alpha-\delta-\varepsilon)}\,\mathrm{d}\xi$
$\lesssim n^{-d}\big\|\widehat{Q}\big\|_{\infty}\|\theta^{n}_{r}\|_{2}^{2},$